# Optimistic Verifiable Training by Controlling Hardware Nondeterminism

### 14 Mar 2024 v1

@@ Abstract
The increasing compute demands of AI systems has led to the emergence of services that train models on behalf of clients lacking necessary resources. However, ensuring correctness of training and guarding against potential training-time attacks, such as data poisoning, poses challenges. Existing works on verifiable training largely fall into two classes: proof-based systems, which struggle to scale due to requiring cryptographic techniques, and “optimistic” methods that consider a trusted third-party auditor who replicates the training process. A key challenge with the latter is that hardware nondeterminism between GPU types during training prevents an auditor from replicating the training process exactly, and such schemes are therefore non-robust. We propose a method that combines training in a higher precision than the target model, rounding after intermediate computation steps, and storing rounding decisions based on an adaptive thresholding procedure, to successfully control for nondeterminism. Across three different NVIDIA GPUs (A40, Titan XP, RTX 2080 Ti), we achieve exact training replication at FP32 precision for both full-training and fine-tuning of ResNet-50 (23M) and GPT-2 (117M) models. Our verifiable training scheme significantly decreases the storage and time costs compared to proof-based systems.
@@ 1 Introduction
We are currently in the “large-scale era” of machine learning (ML), where the exciting capabilities of modern AI systems have required a dramatic increase in training compute needs (Sevilla et al., 2022). In turn, several model training services (e.g. Replicate, OpenAI’s Finetuning API, Together AI, Amazon Sagemaker, MosaicML Training) have been created to support clients who lack the resources to train a model themselves. However, these services require clients to place a significant degree of trust in them to train the model correctly, without any training-time attacks such as data poisoning or undetectable backdoors (Wan et al., 2023; Goldwasser et al., 2022). How can we help a client, who is likely just one individual or a small company, hold the service provider accountable?
@@
Consider an education start-up that wishes to finetune the Llama-70b language model (70B parameters) on their own curated dataset to support student learning. This task requires significant resources, and the company might even lack the necessary expertise. Instead, they might choose to pay a $\mathrm{trainer}$ with vast computing resources to perform the training task (Figure 1). However, what if the $\mathrm{trainer}$ includes data points that spread misinformation, or introduces backdoors that advance a political agenda only for specific prompts? If the client starts to notice suspicious model behavior, is there any action they can take? We study this problem of verifiable training, or ensuring that the training of an ML model was performed correctly.
@@
One possibility is for the the $\mathrm{trainer}$ to provide the client a proof that they did indeed train according to specification. However, proof-based systems require cryptographic techniques that struggle to scale to complexity of real-world ML systems. For instance, recent work based on zero-knowledge proof systems for verifiable inference, a much simpler task than training, requires more than 8 minutes to generate proofs for only 20 images (Liu et al., 2021)! Thus, practical proof-based methods for verifiable training have only been implemented for simple tasks such as logistic and linear regression (Garg et al., 2023; Ames et al., 2022).
@@
An alternative “optimistic” approach is to consider a trusted third-pary $\mathrm{auditor}$, such as a non-profit organization, that has sufficient compute to perform the training task, even if not at the bandwidth of a service provider (Figure 1). When a $\mathrm{client}$ suspects foul play, they can ask the $\mathrm{auditor}$ to challenge the $\mathrm{trainer}$ by training the model using their own compute, and demonstrate that the $\mathrm{trainer}$ did not train correctly. This protocol can be efficiently carried about using techniques from the verifiable computation literature, such as the well-known “verification game” from Teutsch & Reitwießner (2019), which proposes a binary-search procedure to identify the exact intermediate computation steps (e.g. training epochs) that two parties diverge on.
@@
Unfortunately, the issue with such “optimistic” approaches is nondeterminism during model training – two models trained on different GPU types, even with same data ordering and random seed, can learn different weights (Figure 2). The $\mathrm{auditor}$ cannot simply compare their model weights with the $\mathrm{trainer}$’s, and recent work has shown that protocols based on comparing model weights, such as Jia et al. (2021)’s “proof of learning,” are non-robust and can be easily forged due to errors from nondeterminism (Thudi et al., 2022; Fang et al., 2023).
@@
We address this limitation by asking: can the $\mathrm{trainer}$ provide additional information to the $\mathrm{auditor}$ that removes the effects of hardware nondeterminism? Our key insight is that hardware nondeterminism occurs due to the accumulation of errors from floating point operations, and so if we choose to perform computation steps in a higher precision (e.g. FP32) than the target precision of the model weights (e.g. FP16), and periodically round back to the target precision, such errors can be “absorbed” by the lower-precision bits. However, because of values that straddle the rounding boundary, it is still possible for the $\mathrm{trainer}$ and $\mathrm{auditor}$ to round in different directions and diverge. If the $\mathrm{trainer}$ also shares rounding directions for every intermediate computation with the $\mathrm{auditor}$ (which requires far fewer bits than sharing computation outputs), the $\mathrm{auditor}$ can replicate the exact training procedure and eliminate errors due to hardware nondeterminism.
@@
We use this strategy to adapt the verification game described by Teutsch & Reitwießner (2019) for verifiable training. The game’s efficiency lies in our ability to store hashes of model checkpoints in a Merkle tree instead of the entire model weights due to eliminating nondeterminism error (Merkle, 1988). To determine if training was performed correctly, the $\mathrm{auditor}$ only needs to compare the root hash of their Merkle tree with that $\mathrm{trainer}$’s – if they do not match, the two parties can enter an interactive binary search procedure to identify the exact training steps of the dispute. We then show this verifiable training scheme can scale to tasks such as full training of ResNet-50 (23M parameters) and finetuning of GPT-2 (117M parameters), significantly outperforming existing methods with respect to both time and storage cost, while eliminating statistical error due to non-determinism. For example, the proposal in prior work Jia et al. (2021) would require $\mathbf{>140\times}$ more storage cost than our method by saving and comparing model weights at every step in order to achieve the lowest possible (yet still non-zero) statistical error.
@@
Concretely, our contributions include: (1) A novel verifiable training scheme based on the verification game from Teutsch & Reitwießner (2019), which stores model weights in a Merkle tree for efficient comparison between a $\mathrm{trainer}$ and $\mathrm{auditor}$; (2) A method for two parties, training the same model on different GPU types, to achieve identical results based on sharing rounding decisions; (3) Experiments showing the ability of our approach to scale to large models such as ResNet-50 and GPT-2 between three different NVIDIA GPU architectures (A40, Titan XP, RTX 2080 Ti); (4) Methods to reduce the storage cost of our approach via efficient encoding of rounding logs and an adaptive threshold mechanism to reduce the amount of rounding decisions logged; and, (5) Comparisons with existing methods, including proof-based systems, that highlight the improved storage and time efficiency of our method.  [^1] Our method is implemented entirely within the pytorch framework, and is available at https://github.com/meghabyte/verifiable-training.
@@ 2 Related Works
Verifiable ML: Training and running inference on ML models in trusted environments has been an exciting direction explored by many researchers. One line of work consists of proof-based systems, where a proof of correctness (for a desired specification) is provided using cryptographic techniques such as succinct zero-knowledge non-interactive arguments (zk-SNARKs) (Micali, 1994; Lee et al., 2020; Liu et al., 2021; Garg et al., 2023; Kang et al., 2022). The majority of these works focus on the easier problem of verifiable inference, or proving that a prediction indeed came from a given model. However, such methods require approximations for non-linear functions such as softmax and ReLU activations, and require proof times on the order of minutes for a single data input (Liu et al., 2021). Subsequently, even the most recent proof-based systems for verifiable training suffer extreme latency, such as 22 minutes for training VGG-11 on one batch of 16 data inputs (Abbaszadeh et al., 2024), and efficient protocols have only been developed for simple models such as linear and logistic regression, for which we believe it is less likely for a client to delegate out training in the first place (Garg et al., 2023; Ames et al., 2022).
@@
An alternative direction is to consider hardware-based solutions, such as training models in trusted execution environments (TEEs). Such approaches would only require a 3rd-party $\mathrm{auditor}$ to review the source code and training data that is used by the $\mathrm{trainer}$ and attested for by the TEE, and not perform expensive model training themselves as in our approach. Unfortunately, TEEs such as NVIDIA’s recently announced H100 “Confidential GPU” not only incur a high performance penalty for model training due to the I/O operations required for encrypting training data (Dhanuskodi et al., 2023), but also run the risk of losing all security guarantees if any attacker extracts the attestation key, as has been successfully demonstrated by several previous works (Nilsson et al., 2020; Bulck et al., 2018).
@@
Most similar to our approach are proof-of-learning protocols, which also consider a trust third-party that compares checkpointing during the course of training with the original training sequence (Jia et al., 2021). However, such methods not only incur high storage cost by requiring model weights to be stored frequently, but are subject to statistical error from training nondeterminism. Several works have shown that proof-of-learning protocols can be spoofed and fail to verify correctness in several important contexts, including training tasks such as machine unlearning, where the $\mathrm{trainer}$ needs to produce a new model that “forgets” specific training data due to privacy legislation (Fang et al., 2023; Kong et al., 2023; Thudi et al., 2022). Although Choi et al. (2023) recently proposed a verification procedure that is immune to several known proof-of-learning attacks, their method is not only limited to supervised learning algorithms, but also based on an assumption that models temporarily overfit data during training, which may not hold true for all training tasks. Overall, these works argue that the main limitation for provably robust protocols is error due to training nondeterminism, motivating our work.
@@ 3 Set-Up: The Verification Game
Our method for verifiable training is based on the interactive verification game proposed by Teutsch & Reitwießner (2019) in the context of blockchains. The basic idea of the verification game is to resolve a dispute between a challenger, in our case the $\mathrm{auditor}$, and a solver, in our case the $\mathrm{trainer}$, for an expensive computation (e.g. model training). In order for the $\mathrm{auditor}$ to take any meaningful action (e.g. pursue legal action), they need to prove the exact source of the dispute (e.g. training time-step where an attack occurred). If we can save model weights at different time steps $t=1...T$ into a compact data structure such as a Merkle tree, then identifying the source of disagreement can be run efficiently via binary search (Merkle, 1988). More formally, the verification game consists of the following parties:
@@
  • $\mathrm{client}$, who receives a model from the $\mathrm{trainer}$ and is concerned about its correctness. They are unable to perform the training task themselves, and is willing to share their data with both $\mathrm{trainer}$ and $\mathrm{auditor}$.
  • $\mathrm{auditor}$, who officially challenges the $\mathrm{trainer}$ on behalf of a client. This is a trusted 3rd-party that also has sufficient resources to perform the same training task as the $\mathrm{trainer}$, but does not provide such training as a service. While the $\mathrm{auditor}$ could be trusted regulatory organization, our work opens up the possibility of the $\mathrm{trainer}$ relying on another provider for auditing, leveraging the rapid increase in model training services.
@@
When the $\mathrm{trainer}$ is approached by an $\mathrm{auditor}$, they would need to share training parameters, model architecture, and randomness, as shown in Figure 1. The $\mathrm{auditor}$ would then replicate the training process, storing model weights at the same interval $k$ as the trainer did in a Merkle tree, which is a tree where every leaf node is a hash of the model weights and every non-leaf node is a hash of its children nodes. The main loop of the verification game starts when both parties have the root hash of their respective Merkle trees – if training was performed correctly, then the $\mathrm{trainer}$’s root should match the $\mathrm{auditor}$’s. Otherwise, a binary search procedure is performed, where the $\mathrm{auditor}$ iteratively descends the data structure until it identifies two consecutive leaf nodes, $i$ and $i+1$, where the hash at $i$ matches that of the $\mathrm{trainer}$, but not $i+1$, thereby identifying the source of the dispute.
Overall, this interaction requires cooperation with the $\mathrm{trainer}$ – for example, if a $\mathrm{trainer}$ refuses to share the value at a certain node of their Merkle tree within a given time frame, they can be considered to have failed the audit. Additionally, we emphasize that the $\mathrm{trainer}$ and $\mathrm{auditor}$ can only use an efficient Merkle tree structure to store model weights, requiring far less storage than prior work, if correct training produces identical weights (and therefore identical hash values). Unfortunately, training nondeterminism between GPUs leads to weight divergence, which is why we seek to control it.
@@ 4 The Nondeterminism Challenge
Although there exist user-side controls for deterministic operations within a single GPU architecture (e.g. NVIDIA H100), these controls do not prevent nondeterminism between GPU architectures (e.g. NVIDIA H100 and V100), where trained models can have similar aggregate performance (e.g. accuracy) yet yield very different predictions (Figure 2) (Crane, 2018a; NVIDIA, 2022). We now briefly discuss the source of nondeterminism between GPU types.
Floating Point Arithmetic: Computers represent real values using integer and floating-point representations, typically using the IEEE 754 standard (Figure 3). There is a tradeoff between the approximation fidelity and the # of bits used to represent the real values. The chosen precision controls the representable numerical range (e.g. 32-bit floating-point values can represent values between $381.17549435e-$ 38 and $383.40282347e+$ 38). Because computers round to representable floating point values, changing the order in which floating point numbers are accumulated can change the resulting sum (Kahan, 1965; Whitehead & Fit-Florea, 2011).
Example: Floating Point Addition Consider three values $a=0.1$, $b=-0.1$ and $c=0.2$ in FP32. The following two orders of summation — $a+c+b$ vs. $a+b+c$ yield different results — $0.200000017881$ vs. $0.200000002980$.
@@
Over the course of the many operations involved in training, rounding can lead to large differences in end results.
@@
Parallel Computation: In GPUs a single operation (called a kernel) is executed by 1000s of threads in parallel. In NVIDIA terminology, GPUs contain a set of streaming multiprocessors (SMs), which run the thread blocks required for the kernel. At the hardware level, the blocks are divided into warps, or groups of threads, that are assigned to the available cores. Applications partition arithmetic workloads (e.g. batch matrix multiplies), differently depending on the GPU architecture to achieve high performance (NVIDIA, 2022), thus changing the order of floating point operations. Different GPUs have different # and sizes of compute units — for instance, the number and tile dimensions of tensor core units (specialized matrix multiplication units on the GPU) across NVIDIA Volta and Turing vs. Ampere and Hopper architectures varies (Sun et al., 2022).
@@
Non Determinism To compute primitives such as GEMMs ($D=AB+C$), the workhorse of machine learning, GPUs split the work of computing the tiles of $D$ across a thread block (NVIDIA, 2023). The results across threads are accumulated. Due to the varied parallelization and data access patterns across architectures, and the fact that floating point addition is not associative, results across GPU architectures are nondeterministic (Crane, 2018a; NVIDIA, 2022).
Any exact verifiable training method, that does not introduce statistical error, would need to control these sources of non-determinism across GPU types.

### 16 Mar 2024 v2

@@ Abstract
The increasing compute demands of AI systems has led to the emergence of services that train models on behalf of clients lacking necessary resources. However, ensuring correctness of training and guarding against potential training-time attacks, such as data poisoning and trapdoors, poses challenges. Existing works on verifiable training largely fall into two classes: proof-based systems, which can be difficult to scale, and “optimistic” methods that consider a trusted third-party auditor who replicates the training process. A key challenge with the latter is that hardware nondeterminism between GPU types during training prevents an auditor from replicating the training process exactly, and such schemes are therefore non-robust. We propose a method that combines training in a higher precision than the target model, rounding after intermediate computation steps, and storing rounding decisions based on an adaptive thresholding procedure, to successfully control for nondeterminism. Across three different NVIDIA GPUs (A40, Titan XP, RTX 2080 Ti), we achieve exact training replication at FP32 precision for both full-training and fine-tuning of ResNet-50 (23M) and GPT-2 (117M) models. Our verifiable training scheme significantly decreases the storage and time costs compared to proof-based systems.
@@ 1 Introduction
We are currently in the “large-scale era” of machine learning (ML), where the exciting capabilities of modern AI systems have required a dramatic increase in training compute needs (Sevilla et al., 2022). In turn, several model training services, such as Replicate, OpenAI’s Finetuning API, Together AI, Amazon Sagemaker, MosaicML Training, and Gensyn, have been created to support clients who lack the resources to train a model themselves. However, these services require clients to place a significant degree of trust in them to train the model correctly, without introducing a training-time attack such as data poisoning or undetectable backdoors (Wan et al., 2023; Goldwasser et al., 2022). How can we help a client, such as an individual or a small company, hold the service provider accountable in case of misbehavior during training?
@@
Consider an education start-up that wishes to finetune the Llama-70b language model (70B parameters) on their own curated dataset to support student learning. This task requires significant resources, and the company might even lack the necessary expertise. Instead, they might choose to pay a $\mathrm{trainer}$ with vast computing resources to perform the training task (Figure 1). However, what if the $\mathrm{trainer}$ adds data points that spread misinformation, or introduces a backdoor that advances a political agenda for specific prompts? If the client starts to notice suspicious model behavior, is there any action they can take? We study this problem of verifiable training, or ensuring that the training of an ML model was performed correctly.
@@
One possibility is for the $\mathrm{trainer}$ to provide the client with a cryptographic proof that the model was trained according to the specification. However, proof-based systems require cryptographic techniques that can be difficult to scale to the complexity of real-world ML systems. For instance, recent work based on zero-knowledge proof systems for verifiable inference, a much simpler task than training, requires more than 8 minutes to generate proofs for only 20 images (Liu et al., 2021). Thus, practical proof-based methods for verifiable training have only been implemented for simple tasks such as logistic and linear regression (Garg et al., 2023; Ames et al., 2022).
@@
An alternative “optimistic” approach is to consider a trusted third-pary $\mathrm{auditor}$, such as a non-profit organization, that has sufficient computing resources to perform the training task, even if not at the bandwidth of a service provider (Figure 1). When a $\mathrm{client}$ suspects foul play, they can ask the $\mathrm{auditor}$ to challenge the $\mathrm{trainer}$ by training the model using the $\mathrm{auditor}$’s own compute, and demonstrate that the $\mathrm{trainer}$ did not train correctly. This protocol can be efficiently carried out using techniques from the literature on verifiable computing, such as the “verification game” method of Teutsch & Reitwießner (2019). This method uses an interactive binary-search procedure to identify the exact intermediate computation step (e.g. training epoch) where the two parties diverged.
@@
Unfortunately, the issue with such “optimistic” approaches is nondeterminism during model training: two models trained on different GPU types, even with same data order and random seed, can learn different weights (Figure 2). The $\mathrm{auditor}$ cannot simply compare their model weights with the $\mathrm{trainer}$’s, and recent work has shown that protocols based on comparing model weights, such as Jia et al. (2021)’s “proof of learning,” are not robust and can be forged due to errors from nondeterminism (Thudi et al., 2022; Fang et al., 2023).
@@
In this paper we address this limitation by asking: can the $\mathrm{trainer}$ provide additional information to the $\mathrm{auditor}$ that removes the effects of hardware nondeterminism? Our starting point is the observation that hardware nondeterminism occurs due to the accumulation of errors from floating point operations. For example, a matrix-vector multiply often results in different floating point values on different GPUs. A natural approach is to perform training using a higher precision (e.g. FP32) than the target precision of the model weights (e.g. FP16), and periodically round back to the target precision. The hope is that all floating point errors will be confined to the higher precision bits, so that rounding will “clean” the computed values and ensure that the trainer and auditor stay in sync. However, this does not work. The reason is that the computed values occasionally straddle the rounding boundary, causing the $\mathrm{trainer}$ to round up while the $\mathrm{auditor}$ rounds down, quickly causing them to diverge. Alternatively, we propose a solution where the $\mathrm{trainer}$ records the rounding direction for every intermediate computation so that $\mathrm{auditor}$ can stay in sync with the $\mathrm{trainer}$. However, as this requires the $\mathrm{trainer}$ to record and output a large number of bits, we also show how to reduce the amount of data needed to eliminate errors due to hardware nondeterminism, as needed for correct auditing.
@@
We use this strategy to adapt the verification game described by Teutsch & Reitwießner (2019) for verifiable training. The game’s efficiency lies in our ability to store hashes of model checkpoints in a Merkle tree (Merkle, 1988). To determine if training was performed according to the specification, the $\mathrm{auditor}$ needs to reconstruct the Merkle tree and compare the resulting Merkle root hash with the Merkle root hash provided by the $\mathrm{trainer}$’s – if they do not match, the two parties can enter an interactive binary search procedure to identify the exact training step of the dispute. The purpose of the binary search game is to hold both parties accountable: an $\mathrm{auditor}$ should not be able to simply claim that a model was improperly trained, but convince a third-party (e.g., the public, or a $\mathrm{judge}$) by showing at what point during training the $\mathrm{trainer}$ misbehaved. We show this verifiable training scheme can scale to tasks such as full training of ResNet-50 (23M parameters) and finetuning of GPT-2 (117M parameters), significantly outperforming existing methods with respect to both time and storage cost, while eliminating statistical error due to non-determinism. For example, the proposal in prior work Jia et al. (2021) would require $\mathbf{>140\times}$ more storage cost than our method by saving and comparing model weights at every step in order to achieve the lowest possible (yet still non-zero) statistical error.
@@
Concretely, our contributions include: (1) A method for two parties, training the same model on different GPU types, to achieve identical results based on sharing rounding decisions; (2) A verifiable training scheme based on the verification game from Teutsch & Reitwießner (2019), which stores model weights in a Merkle tree for efficient comparison between a $\mathrm{trainer}$ and $\mathrm{auditor}$; (3) Experiments showing the ability of our approach to scale to large models such as ResNet-50 and GPT-2 between three different NVIDIA GPU architectures (A40, Titan XP, RTX 2080 Ti); (4) Methods to reduce the storage cost of our approach via efficient encoding of rounding logs and an adaptive threshold mechanism to reduce the amount of rounding decisions logged; and (5) Comparisons with existing methods, including proof-based systems, that highlight the improved storage and time efficiency of our method.  [^1] Our method is implemented entirely within the pytorch framework, and is available at https://github.com/meghabyte/verifiable-training.
@@ 2 Related Works
Verifiable ML: Training and running inference on ML models in trusted environments has been an exciting direction explored by many researchers. One line of work consists of proof-based systems, where a proof of correctness (for a desired specification) is provided using cryptographic techniques such as succinct non-interactive arguments (SNARKs) (Micali, 1994; Bitansky et al., 2012; Lee et al., 2020; Liu et al., 2021; Garg et al., 2023; Kang et al., 2022). The majority of these works focus on the easier problem of verifiable inference, or proving that a prediction indeed came from a given model. However, such methods require approximations for non-linear functions such as softmax and ReLU activations, and require proof times on the order of minutes for a single data input (Liu et al., 2021). Subsequently, even the most recent proof-based systems for verifiable training suffer extreme latency, such as 22 minutes for training VGG-11 on one batch of 16 data inputs (Abbaszadeh et al., 2024), and efficient protocols have only been developed for simple models such as linear and logistic regression, for which it is less likely for a client to delegate out training in the first place (Garg et al., 2023; Ames et al., 2022).
@@
An alternative direction is to consider hardware-based solutions, such as training models in a trusted execution environment (TEE). Such approaches would only require a 3rd-party $\mathrm{auditor}$ to review the source code and training data that is used by the $\mathrm{trainer}$ and attested for by the TEE, and not perform expensive model training themselves as in our approach. However, TEEs such as NVIDIA’s H100 “Confidential GPU” incur a performance penalty for model training due to the cost of running inside a TEE (Dhanuskodi et al., 2023). Moreover, clients run the risk of losing all security guarantees if an attacker can extract the attestation key from even one GPU, as has been successfully demonstrated in several works (Nilsson et al., 2020; Bulck et al., 2018).
@@
Most similar to our approach are proof-of-learning protocols, which also consider a trusted 3rd party that compares checkpointing during the course of training with the original training sequence (Jia et al., 2021). However, such methods not only incur high storage cost by requiring model weights to be stored frequently, but are also subject to errors from training nondeterminism, and cannot provide a robust guarantee. Several works have shown that proof-of-learning protocols can be spoofed and fail to verify correctness in several important contexts, including training tasks such as machine unlearning, where the $\mathrm{trainer}$ needs to produce a new model that “forgets” specific training data due to privacy legislation (Fang et al., 2023; Kong et al., 2023; Thudi et al., 2022). Although Choi et al. (2023) recently proposed a verification procedure that is immune to several known proof-of-learning attacks, their method is not only limited to supervised learning algorithms, but also based on an assumption that models temporarily overfit data during training, which may not hold true for all training tasks. Overall, these works argue that the main limitation for provably robust protocols is error due to training nondeterminism. This provides further motivation for our work.
@@ 3 Set-Up: The Verification Game
Our method for verifiable training is based on the interactive verification game proposed by Teutsch & Reitwießner (2019) in the context of blockchains. The basic idea of the verification game is to resolve a dispute between a challenger, in our case the $\mathrm{auditor}$, and a solver, in our case the $\mathrm{trainer}$, for an expensive computation (e.g. model training). In order for the $\mathrm{auditor}$ to take any meaningful action (e.g. pursue legal action), they need to prove the exact source of the dispute (e.g. training time-step where an attack occurred). If we can save model weights at different time steps into a compact data structure such as a Merkle tree, then identifying the source of disagreement can be done efficiently using binary search (Merkle, 1988). More precisely, the verification game consists of the following parties:
@@
  • $\mathrm{client}$, who receives a model from the $\mathrm{trainer}$ and is concerned about its correctness. They are unable to perform the training task themselves, and are willing to share their data with both $\mathrm{trainer}$ and $\mathrm{auditor}$.
  • $\mathrm{auditor}$, who officially challenges the $\mathrm{trainer}$ on behalf of a client. This is a trusted 3rd-party that also has sufficient resources to perform the same training task as the $\mathrm{trainer}$, but does not provide such training as a service. If needed, the client can choose several auditors to audit the model that was generated by the $\mathrm{trainer}$. While an $\mathrm{auditor}$ could be a trusted regulatory organization, it could also be some other provider of training services.
@@
When the $\mathrm{trainer}$ is approached by an $\mathrm{auditor}$, they would need to share training parameters, model architecture, and randomness, as shown in Figure 1. The $\mathrm{auditor}$ would then replicate the training process, storing model weights in a Merkle tree at the same checkpoints as the trainer. Recall that a Merkle tree is a tree where every leaf node is a hash of the model weights and every non-leaf node is a hash of its children nodes. The main loop of the verification game starts when both parties have the root hash of their respective Merkle trees. If training was performed correctly, then the $\mathrm{trainer}$’s root should match the $\mathrm{auditor}$’s. Otherwise, a binary search procedure is performed, where the $\mathrm{auditor}$ iteratively descends the Merkle tree until it identifies two consecutive leaf nodes, $i$ and $i+1$, where the hash at $i$ matches that of the $\mathrm{trainer}$, but the hash at leaf $i+1$ does not. This identifies the point in the computation that caused the dispute.
This interactive verification game requires the cooperation of the $\mathrm{trainer}$. If the $\mathrm{trainer}$ refuses to share the value at a certain node of their Merkle tree within a given time frame, they can be considered to have failed the audit. Additionally, we emphasize that the $\mathrm{trainer}$ and $\mathrm{auditor}$ use a Merkle tree to store model weights, requiring far less storage than prior work, if correct training produces identical weights (and therefore identical hash values).
The problem is that training nondeterminism between GPUs leads to weight divergence between the auditor and trainer, and causes the verification game outlined above to always fail. This why we seek to prevent divergence in training.
@@ 4 The Nondeterminism Challenge
Although there are user-side controls for forcing deterministic operations within a single GPU architecture (e.g. NVIDIA H100), these controls do not prevent nondeterminism between GPU architectures (e.g. NVIDIA H100 and V100), where trained models can have similar aggregate performance (e.g. accuracy) yet yield very different predictions, as shown in Figure 2 (Crane, 2018a; NVIDIA, 2022). We now discuss the source of nondeterminism between GPU types.
Floating Point Arithmetic: Computers represent real values using integer and floating-point representations, typically using the IEEE 754 standard (Figure 3). There is a tradeoff between the approximation fidelity and the number of bits used to represent the real values. The chosen precision controls the representable numerical range (e.g. 32-bit floating-point values can represent values between $381.17549435e-$ 38 and $383.40282347e+$ 38). Because computers round to representable floating point values, changing the order in which floating point numbers are accumulated can change the resulting sum (Kahan, 1965; Whitehead & Fit-Florea, 2011).
Example: Floating Point Addition Consider three values $a=0.1$, $b=-0.1$ and $c=0.2$ in FP32. The following two orders of summation, $a+c+b$ vs. $a+b+c$, yield different results: $0.200000017881$ vs. $0.200000002980$.
@@
Over the course of the many operations during training, this rounding to a representable floating point value can lead to a large difference in the end result between the $\mathrm{trainer}$ and $\mathrm{auditor}$.
@@
Parallel Computation: In a GPU, a single operation (called a kernel) is executed by thousands of threads in parallel. In NVIDIA terminology, GPUs contain a set of streaming multiprocessors (SMs), which run the thread blocks required for the kernel. At the hardware level, the blocks are divided into warps, or groups of threads, that are assigned to the available cores. Applications partition arithmetic workloads (e.g. batch matrix multiplies), differently depending on the GPU architecture to achieve high performance (NVIDIA, 2022), thus changing the order of floating point operations. Different GPUs have a different number and size of compute units. For instance, the number and tile dimensions of tensor core units (specialized matrix multiplication units on the GPU) on NVIDIA Volta and Turing vs. Ampere and Hopper architectures is different (Sun et al., 2022). As a result, a kernel executed on Volta will result in a different value than a kernel executed on Hopper. Other work uses nondeterministic execution patterns, allowing a worker to perform redundant work when it is faster than waiting on sequentially dependent outputs from a slower parallel processor, to improve efficiency (Merrill & Garland, 2016).
@@
Non Determinism To compute primitives such as GEMMs ($D=A\cdot B+C$), the workhorse of machine learning, GPUs split the work of computing the tiles of $D$ across a thread block (NVIDIA, 2023). The results across threads are accumulated. Due to the varied parallelization and data access patterns across architectures, and the fact that floating point addition is not associative, results across GPU architectures are nondeterministic (Crane, 2018a; NVIDIA, 2022).
Any exact and robust verifiable training method would need to control these sources of non-determinism across GPUs.